Project

General

Profile

Commissioning and Ops Scheduling meeting : 10/11/2015

To be held in CR10 in the ATLAS Building @ 15:00 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Introduction (Steve B)
  • DC Report (Paul H)
  • Commissioning Tasks and Schedule for Cycle 2015/03 and beyond : Chaired by Paul H
    • Current known commissioning tasks
      • Focus coil flip mode commissioning (Steve W)
      • Spectrometer Solenoid work (this needs specification)
      • Absorber work (Steve W)
      • Diffuser work (?)
      • Others ?
    • Data-taking possibilities (Chris R)
    • Maintenance plans
      • Reboot of the vacuum controls (P. Warburton)
      • Decay solenoid power supply repair (?)
      • Cooling water work (Colin)
      • Compressed air transducer installation (Henry)
  • Change of meeting time
  • AOB

Minutes

Present: Holger, Victoria, Craig, Celeste, Paul, Henry, Steve W, Mark Tucker, Steve B, Andg G, Yordan, John G, Ed Overton,
Mike Courthold, John Nugent, Chris R

Focus Coil Commissioning : Thursday 12th / Friday 13th November
  • Close and search Hall on Wednesday afternoon
  • Start ramping FC to 225 A on Thursday morning, then ramp down to 200 A and do a 24 hour soak test
Spectrometer Solenoid :
  • will do some work on SS power supplies and vacuum controls on Thursday 12/Friday 13 and into next week
Absorber:
  • Leak testing to take place Wednesday 11th to Monday 16th
  • Possibly cool down Thursday 12th to Monday 16th
Diffuser: There is the possibility to change the diffuser in the week Monday 16th - Friday 20th.
  • Decision to be made on Wednesday 11th
  • If so, TOF1 will be moved and PRY opened on Monday 16th
  • Top hat will then be removed on Monday afternoon and sent to Oxford
  • Reinstallation may take the rest of the week (allowing for contingency). This is the major work next week, if it goes ahead.

Hall: John wants to change the fluorescent tubes for LEDs at some point, but this is not urgent.

Control room (added during discussion after meeting): Durga and the online group want to take a couple of days to test control room systems.
This will take place towards the end of next week (Wed 18th - Friday 20th November).

Data-taking possibilities: Chris presented a framework for running at the end of the December. Data-taking has
the goals of
  • FC alignment : requires the FC to be operating at low current
  • Multiple Coulomb Scattering measurement with straight tracks : requires LH2 absorber
  • Precision measurement of input emittance : requires E-C-E coils on from SSU

No decision made as yet as input required from magnet review as well as better understanding of absorber schedule.
Nevertheless, should plan for some running so will probably open shifts from December 5th to December 18th.

  • Plan for an expert-led run for Tuesday 24th or Thursday 26th November

Maintenance:

  • Decay solenoid being looked at this afternoon (Tuesday 10th) and a decision made on what to do. DS could be off until January.
  • Work on second cooling water pump controller delayed due to concerns about cryomechs. Will look at flushing conventional
    magnet cooling system on Wednesday 25th (ISIS maintenance day) first.
  • Extended debate about whether the compressed air transducer Henry wants to install is necessary. Henry will discuss with Mike and John.
  • Reboot of vacuum controls will take place on the morning of Wednesday 11th.

AOB: Henry suggested that we sort out scheduling of 800 MeV activation run with Dean ASAP. Will bring this up at MICE/ISIS meeting on
Wednesday 11th

Date of next meeting: Due to possible clash with MEMO/MIPO meetings on Tuesday afternoons, this co-ordination meeting was
shifter half an hour back. It will now take place on Tuesday's at 15:30 BST.